The manager of the machine-made kretek kretek cigarette company RAENAK found that 12% of its production was below production standards (poor). One day, the manager received a report that on that day 17% of substandard cigarette products were obtained from a random sample of 1000 cigarettes. Test whether the manufacturing defects (17%) are random or permanent (machines need to be repaired)! Use a= 0.05!

1
Expert's answer
2022-06-06T11:16:11-0400

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\le0.12"

"H_a:p>0.12"

This corresponds to a right-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05

," and the critical value for a right-tailed test is "z_c =1.6449."

The rejection region for this right-tailed test is "R = \{z: z > 1.6449\}."

The z-statistic is computed as follows:




"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.17-0.12}{\sqrt{\dfrac{0.12(1-0.12)}{1000}}}=4.8656"

Since it is observed that "z = 4.8656>1.6449= z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p=P(Z>4.8656)=0.000001," and since "p=0.000001<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p" is greater than 0.12, at the "\alpha = 0.05" significance level.